# Kennedy-Tasaki transformation and non-invertible symmetry in lattice models beyond one dimension

### v1

@@ Abstract
We give an explicit operator representation (via a sequential circuit and projection to symmetry subspaces) of Kramers-Wannier duality transformation in higher-dimensional subsystem symmetric models generalizing the construction in the 1D transverse-field Ising model. Using the Kramers-Wannier duality operator, we also construct the Kennedy-Tasaki transformation that maps subsystem symmetry-protected topological phases to spontaneous subsystem symmetry breaking phases, where the symmetry group for the former is either $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ or $\mathbb{Z}_{2}$. This generalizes the recently proposed picture of one-dimensional Kennedy-Tasaki transformation as a composition of manipulations involving gauging and stacking symmetry-protected topological phases to higher dimensions.
@@ I Introduction
Conventional spontaneous symmetry breaking (SSB) phases, in contrast to SPT phases, can be described by local order parameters, with the famous example of the 2D classical Ising model [11, 12], as well as its corresponding 1D quantum Ising model in a transverse field. In these models, the so-called Kramers-Wannier (KW) duality [13, 14] maps the model to one on the dual lattice between high-temperature (or high-field, i.e., disordered) and low-temperature (or low-field, i.e., ordered) phases. There is a recent re-emergence of interest in the KW duality of the 1D transverse-field Ising model, which is regarded as a transformation on the same lattice, i.e., in the same Hilbert space, exchanging the Ising interaction and transverse-field terms. In addition to the $Z_{2}$ spin-flip symmetry, at the critical point, an additional symmetry is hence given by the KW duality. Recently, an explicit expression for this symmetry action in terms of unitaries and projection was obtained by Refs. [15, 16, 17]. Due to the explicit projection onto the symmetric subspace, this is a non-invertible symmetry. Furthermore, this symmetry action squares to a lattice translation by one site times the projection [16]. Seiberg and Shao also show that this non-invertible symmetry comes from gauging the fermion parity of free Majorana fermions [16]. Majorana translation symmetry emerges as the non-invertible KW duality symmetry after gauging the fermion parity. This non-invertible symmetry fits with the non-invertible duality line in the Ising CFT [18, 19, 20, 21].
The Haldane phase, now recognized as an SPT phase, was found by Kennedy and Tasaki (KT) to relate via a non-local transformation to a spontaneous $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry breaking phase [22, 23]. Oshikawa explicitly constructed such a non-local transformation and generalized it to all integer spins [24]. The overall picture that emerges is that the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SPT is mapped to two copies of the symmetry broken phase under the KT transformation [25]. An explicit example given in Ref. [25] is between the spin-1/2 1D cluster-state Hamiltonian, which represents a nontrivial $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SPT phase, and two copies of the transverse-field Ising model, which represent $\mathbb{Z}_{2}$ SSB phases.
@@
In this paper, we generalize the explicit operator construction of KW duality symmetry to higher dimensional hypercubic Ising models that have subsystem line-like symmetry. Such duality maps between the SSB and disordered phases. We also discuss a fermionic dual to these models: Majorana hypercubic models that have subsystem fermion parity symmetry. We gauge the subsystem fermion parity to obtain the hypercubic Ising models. Connecting SPT to SSB phases, we provide an explicit operator representation of the KT transformation in one and higher dimensions. The KT transformation maps the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-symmetric cluster-state model on the hypercubic bipartite lattice to two copies of hypercubic Ising models. We generalize the picture provided by Ref. [29] for KT transformation to that between $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ subsystem SPT (SSPT) to (two copies of) $\mathbb{Z}_{2}$ subsystem SSB (SSSB) phases in higher dimensions. We also discuss about $\mathbb{Z}_{2}$ subsystem symmetric model in two and higher dimensions and their corresponding $\mathbb{Z}_{2}$ subsystem SPT models. Specifically, we provide an explicit operator representation of the KW duality symmetry for the double hypercubic Ising model (DHCIM), and using this, we construct the KT transformation that maps between $\mathbb{Z}_{2}$ SSPT and one copy of DHCIM (which has SSSB).
@@
The remaining structure of this paper is as follows. We give an overview of the results in Ref. [16] in section II, where they obtain an explicit expression for KW duality on a lattice. Using these results, we give an explicit expression for Kennedy-Tasaki transformation in one dimension mapping $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SPT to two copies of the Ising model. We also discuss the composition of KT and KW. In section III, we discuss our results on the explicit KW duality operator for subsystem symmetric models in two and higher dimensions. Section IV deals with KT transformation that maps the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SSPT to two copies of hypercubic Ising models that are in the SSSB phase. We mention the composition of KW and KT in two and higher dimensions in section V. Discussion of KW and KT transformation for $\mathbb{Z}_{2}$ subsystem symmetric model in two and higher dimensions is presented in section VI. Finally, in section VII, we conclude our results and give some future directions. Some discussions about the fermionic duals of the hypercubic Ising models are given in appendix A. We discuss an alternate way of obtaining the KT transformation in appendix B. In appendix C, we discuss $\mathbb{Z}_{N}$ generalization of hypercubic clock models and their non-invertible symmetries. In appendix D, we provide the $\mathbb{Z}_{N}$ generalization of KT transformation that maps $\mathbb{Z}_{N}\times\mathbb{Z}_{N}$ SPT phase to two copies of $\mathbb{Z}_{N}$ symmetry breaking phase in one dimension and discuss a KT transformation that maps $\mathbb{Z}_{N}\times\mathbb{Z}_{N}$ SSPT phase to two copies of $\mathbb{Z}_{N}$ subsystem symmetry breaking phase in two dimensions. Finally we discuss a way to obtain the Hamiltonian of $\mathbb{Z}_{2}$ SSPT phase from the Hamiltonian of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SSPT by breaking the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry to the diagonal subgroup in appendix E.
@@
II Non-invertible symmetry and Kennedy-Tasaki transformation in one dimensional lattice model
@@
In this section, we review non-invertible symmetry and Kennedy-Tasaki transformation in one dimensional lattice models.
@@ II.1.1 $\mathbb{Z}_{2}$ symmetric model
Two dimensional classical Ising model possesses the famous Kramers-Wannier (KW) duality between high temperature and low temperature phases [13]. The duality maps an ordered phase to a disordered phase and vice versa. At the critical temperature, the KW duality is a self-duality and that in turn determines the critical temperature. Two dimensional classical model can be mapped to a one dimensional quantum model by the quantum-classical correspondence [14]. Hence, one can discuss the KW duality in the context of one dimensional quantum lattice model, which is the transverse field Ising model (TFI). Let us consider the Hamiltonian of the TFI model on a ring with $L$ sites with a coupling parameter $\lambda$
@@
Here $X$, $Y$, and $Z$ are the Pauli operators. They are bosonic operators that anticommute with each other and square to identity. The model has a global $\mathbb{Z}_{2}$ symmetry operator $\eta\equiv\prod_{j=1}^{L}X_{j}$ and it commutes with the Hamiltonian $[\bm{H}_{\text{TFI}},\eta]=0$. This model at $\lambda=1$ also has a symmetry that interchanges the Ising and transverse field term. This is the quantum version of the the celebrated KW self-duality at the critical point for the classical model. Recently, an explicit expression for the KW duality symmetry for TFI was obtained by Ref. [15, 16, 17]. The KW duality which we denote by $\mathbf{D}$ can be decomposed into a unitary part and a projection onto the symmetric subspace where $\eta=1$. The unitary part that we denote by $\tilde{\mathbf{D}}$ is an ordered product of operators (specifically, a Clifford circuit)
@@
{tikzpicture}
Figure 1: Quantum circuit representation of the operator $\tilde{\mathbf{D}}$ given in Eq. (2). Time ordering of the circuit is from top to bottom. Top layer denotes the input sites $1$,$2$,…,$L$. The bottom layer represents the output.
@@
From Eq. (5), we easily see that $\mathbf{D}^{2}X_{j}=X_{j+1}\mathbf{D}^{2}$ and $\mathbf{D}^{2}Z_{j}Z_{j+1}=Z_{j+1}Z_{j+2}\mathbf{D}^{2}$ (with the periodic boundary condition $L+1\equiv 1$). Hence, $\mathbf{D}^{2}$ is a product of lattice translation by one site, which we denote by $T$, and the projection $\mathbf{P}$ onto the symmetric subspace,
$\mathbf{D}^{2}=T\mathbf{P},$ (8)
up to an overall phase factor. Otherwords, we could redefine $\mathbf{D}$ with appropriate phase choice to obtain Eq. (8). As a note, translation operation cannot be constructed in constant depth quantum circuit [30]. This is consistent with the fact that $\tilde{\mathbf{D}}$ is a linear-depth quantum circuit. Additionally, we have omitted the phase factor $e^{-2\pi iN/8}$ related to an anomaly in the corresponding fermionic theory [16] , as it does not affect the transformation of operators.
@@ II.1.2 $\mathbb{Z}_{N}$ symmetric model
with periodic identifications $Z_{i+L}\equiv Z_{i}$ and $X_{i+L}\equiv X_{i}$. This model has a global symmetry generated by $\eta=\prod_{i}X_{i}$. Note that $\eta^{2}$,…, $\eta^{N-1}$ are also global symmetries. The model is known to have a critical point describing phase transition at $\lambda=1$ for $N=2,3,4$  [31, 32, 33]. At $\lambda=1$, there is a symmetry under the Kramers-Wannier transformation that interchanges the interaction term and the transverse-field term. This self-dual point is described by the $\frac{SU(2)_{N}}{U(1)}$ coset conformal field theory with central charge $c=\frac{2(N-1)}{N+2}$ [34, 35]. Let us denote the Kramers-Wannier duality symmetry for this model as $\mathbf{D}_{(N)}$. We define
$\begin{split}\mathbf{D}_{(N)}&\equiv\mathbf{P}_{(N)}\tilde{\mathbf{D}}_{(N)}\mathbf{P}_{(N)}\equiv\mathbf{P}_{(N)}\prod_{j=1}^{L-1}\left[\left(\sum_{n=0}^{N-1}c(n)X_{j}^{n}\right)\left(\sum_{n=0}^{N-1}\tilde{c}(n)(Z_{j}^{\dagger}Z_{j+1})^{n}\right)\right]\times\left(\sum_{n=0}^{N-1}c(n)X_{L}^{n}\right)\mathbf{P}_{(N)},\end{split}$ (10)
@@
$\tilde{c}(n)=\omega^{\frac{n(N+n)}{2}},\quad c(n)=\omega^{-\frac{n(N-n)}{2}}.$ (12a)
@@
$\mathbf{D}_{(N)}X_{j}=Z_{j}^{\dagger}Z_{j+1}\mathbf{D}_{(N)},$ (13a)
$\mathbf{D}_{(N)}X_{j}^{\dagger}=Z_{j}Z_{j+1}^{\dagger}\mathbf{D}_{(N)},$ (13b)
$\mathbf{D}_{(N)}Z_{j}^{\dagger}Z_{j+1}=X_{j+1}\mathbf{D}_{(N)},$ (13c)
$\mathbf{D}_{(N)}Z_{j}Z_{j+1}^{\dagger}=X_{j+1}^{\dagger}\mathbf{D}_{(N)}.$ (13d)
Moreover, $\mathbf{D}_{(N)}$ commutes with the Hamiltonian at $\lambda=1$ and $\mathbf{D}_{(N)}^{2}=T_{clock}\mathbf{P}_{(N)}$ after an appropriate phase choice for $\mathbf{D}_{(N)}$. Here $T_{clock}$ is the translation on the lattice by one site.
@@
$\tilde{\mathbf{D}}_{(N)}Z_{i}=\omega^{-\frac{(N+1)}{2}}Z_{1}\prod_{j=1}^{i}X_{j}\tilde{\mathbf{D}}_{(N)},$ (14)
which reduces to Eq. (6) for $N=2$, up to an irrelevant overall minus sign.

### v2

@@ Abstract
We give an explicit operator representation (via a sequential circuit and projection to symmetry subspaces) of Kramers-Wannier duality transformation in higher-dimensional subsystem symmetric models, generalizing the construction in the 1D transverse-field Ising model. Using the Kramers-Wannier duality operator, we also construct the Kennedy-Tasaki transformation that maps subsystem symmetry-protected topological phases to spontaneous subsystem symmetry-breaking phases, where the symmetry group for the former is either $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ or $\mathbb{Z}_{2}$. This also generalizes the recently proposed picture of the one-dimensional Kennedy-Tasaki transformation as a composition of manipulations involving gauging and stacking symmetry-protected topological phases to higher dimensions.
@@ I Introduction
Conventional spontaneous symmetry breaking (SSB) phases, in contrast to SPT phases, can be described by local order parameters, with the famous example of the 2D classical Ising model [11, 12], as well as its corresponding 1D quantum Ising model in a transverse field. In these models, the so-called Kramers-Wannier (KW) duality [13, 14, 15, 16] maps the model to one on the dual lattice between high-temperature (or high-field, i.e., disordered) and low-temperature (or low-field, i.e., ordered) phases. There is a recent re-emergence of interest in the KW duality of the 1D transverse-field Ising model, which is regarded as a transformation on the same lattice, i.e., in the same Hilbert space, exchanging the Ising interaction and transverse-field terms. In addition to the $Z_{2}$ spin-flip symmetry, at the critical point, an additional symmetry is hence given by the KW duality. Recently, an explicit expression for this symmetry action in terms of unitaries and projection was obtained by Refs. [17, 18, 19]. Due to the explicit projection onto the symmetric subspace, this is a non-invertible symmetry. Furthermore, this symmetry action squares to a lattice translation by one site times the projection [18]. Seiberg and Shao also show that this non-invertible symmetry comes from gauging the fermion parity of free Majorana fermions [18]. Majorana translation symmetry emerges as the non-invertible KW duality symmetry after gauging the fermion parity. This non-invertible symmetry fits with the non-invertible duality line in the Ising CFT [20, 21, 22, 23].
The Haldane phase, now recognized as an SPT phase, was found by Kennedy and Tasaki (KT) to relate via a non-local transformation to a spontaneous $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry breaking phase [24, 25]. Oshikawa explicitly constructed such a non-local transformation and generalized it to all integer spins [26]. The overall picture that emerges is that the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SPT is mapped to two copies of the symmetry broken phase under the KT transformation [27]. An explicit example given in Ref. [27] is between the spin-1/2 1D cluster-state Hamiltonian, which represents a nontrivial $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SPT phase, and two copies of the transverse-field Ising model, which represent $\mathbb{Z}_{2}$ SSB phases. Recently, there have also been generalizations of KT transformation to categorical symmetries [28] and in the context of LDPC codes [29].
@@
In this paper, we generalize the explicit operator construction of the KW duality symmetry to higher-dimensional hypercubic Ising models that have subsystem line-like symmetry. Such duality maps between the SSB and disordered phases. We also discuss a fermionic dual to these models: Majorana hypercubic models that have subsystem fermion parity symmetry. We gauge the subsystem fermion parity to obtain the hypercubic Ising models. Connecting SPT to SSB phases, we provide an explicit operator representation of the KT transformation in one and higher dimensions. The KT transformation maps the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-symmetric cluster-state model on the hypercubic bipartite lattice to two copies of hypercubic Ising models. We generalize the picture provided by Ref. [33] for KT transformation to that between $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ subsystem SPT (SSPT) to (two copies of) $\mathbb{Z}_{2}$ subsystem SSB (SSSB) phases in higher dimensions. We also discuss about $\mathbb{Z}_{2}$ subsystem symmetric model in two and higher dimensions and their corresponding $\mathbb{Z}_{2}$ subsystem SPT models. Specifically, we provide an explicit operator representation of the KW duality symmetry for the double hypercubic Ising model (DHCIM), and using this, we construct the KT transformation that maps between $\mathbb{Z}_{2}$ SSPT and one copy of DHCIM (which has SSSB).
@@
The remaining structure of this paper is as follows. We give an overview of the results in Ref. [18] in section II, where they obtain an explicit expression for KW duality on a lattice. Using these results, we give an explicit expression for Kennedy-Tasaki transformation in one dimension mapping $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SPT to two copies of the Ising model. We also discuss the composition of KT and KW. In section III, we discuss our results on the explicit KW duality operator for subsystem symmetric models in two and higher dimensions. Section IV deals with the KT transformation that maps the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SSPT to two copies of hypercubic Ising models that are in the SSSB phase. We mention the composition of KW and KT in two and higher dimensions in section V. The discussion of KW and KT transformation for $\mathbb{Z}_{2}$ subsystem symmetric model in two and higher dimensions is presented in section VI. Finally, in section VII, we conclude our results and give some future directions. Some discussions about the fermionic duals of the hypercubic Ising models are given in the appendix A. We discuss an alternate way to obtain the KT transformation in the appendix B. In appendix C, we discuss $\mathbb{Z}_{N}$ generalization of hypercubic clock models and their non-invertible symmetries. In appendix D, we provide the $\mathbb{Z}_{N}$ generalization of KT transformation that maps $\mathbb{Z}_{N}\times\mathbb{Z}_{N}$ SPT phase to two copies of $\mathbb{Z}_{N}$ symmetry breaking phase in one dimension and discuss a KT transformation that maps $\mathbb{Z}_{N}\times\mathbb{Z}_{N}$ SSPT phase to two copies of $\mathbb{Z}_{N}$ subsystem symmetry breaking phase in two dimensions. We discuss a way to obtain the Hamiltonian of $\mathbb{Z}_{2}$ SSPT phase from the Hamiltonian of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SSPT by breaking the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry to the diagonal subgroup in appendix E. In appendix F, we describe how one can implement the KW and KT transformations in short-depth operations using gates and measurements. We discuss KT transformations between order parameters in Appendix G.
@@
II Non-invertible symmetry and Kennedy-Tasaki transformation in one-dimensional lattice model
@@
In this section, we review non-invertible symmetry and Kennedy-Tasaki transformation in one-dimensional lattice models.
@@ II.1.1 $\mathbb{Z}_{2}$ symmetric model
The two-dimensional classical Ising model possesses the famous Kramers-Wannier (KW) duality between high-temperature and low-temperature phases [13]. The duality maps an ordered phase to a disordered phase and vice versa. At the critical temperature, the KW duality is a self-duality, and that in turn determines the critical temperature. The two-dimensional classical model can be mapped to a one-dimensional quantum model by the quantum-classical correspondence [16]. Hence, one can discuss the KW duality in the context of a one-dimensional quantum lattice model, which is the transverse field Ising model (TFI). Consider the Hamiltonian of the TFI model on a ring with $L$ sites with a coupling parameter $\lambda$
@@
Here, $X$, $Y$, and $Z$ are the Pauli operators. They are bosonic operators that anticommute with each other and square to identity. The model has a global $\mathbb{Z}_{2}$ symmetry operator $\eta\equiv\prod_{j=1}^{L}X_{j}$ and it commutes with the Hamiltonian $[\bm{H}_{\text{TFI}},\eta]=0$. This model at $\lambda=1$ also has a symmetry that interchanges the Ising and transverse field term. This is the quantum version of the celebrated KW self-duality at the critical point for the classical model. Recently, an explicit expression for the KW duality symmetry for TFI was obtained by Ref. [17, 18, 19]. The KW duality which we denote by $\mathbf{D}$ can be decomposed into a unitary part and a projection onto the symmetric subspace where $\eta=1$. The unitary part that we denote by $\tilde{\mathbf{D}}$ is an ordered product of operators (specifically, a Clifford circuit)
@@
Figure 1: Quantum circuit representation of the operator $\tilde{\mathbf{D}}$ given in Eq. (2). The time ordering of the circuit is from top to bottom. The top layer denotes the input sites $1$,$2$,…,$L$. The bottom layer represents the output.
@@
From Eq. (5), we easily see that $\mathbf{D}^{2}X_{j}=X_{j+1}\mathbf{D}^{2}$ and $\mathbf{D}^{2}Z_{j}Z_{j+1}=Z_{j+1}Z_{j+2}\mathbf{D}^{2}$ (with the periodic boundary condition $L+1\equiv 1$). Moreover, it could be easily checked that $\mathbf{D}^{\dagger}\mathbf{D}=\textbf{P}$. Hence, $\mathbf{D}^{2}$ is a product of lattice translation by one site, which we denote by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{T}}$, and the projection $\mathbf{P}$ onto the symmetric subspace,
$\mathbf{D}^{2}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\propto}}\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{T}}\mathbf{P}.$ (8)
The proportionality factor is a phase and is fixed by the action of $\mathbf{D}^{2}$ on the product state |$\ket{+}^{\otimes L}$.
$\mathbf{D}^{2}=e^{\frac{2\pi iL}{4}}\mathbf{T}\mathbf{P}.$ (9)
As a note, translation operation cannot be constructed in constant depth quantum circuit [34]. This is consistent with the fact that $\tilde{\mathbf{D}}$ is a linear-depth quantum circuit. We note that the phase factor $(e^{2\pi iN/8})^{2}$ related to an anomaly in the corresponding fermionic theory [18] , and it does not affect the transformation of operators.
@@ II.1.2 $\mathbb{Z}_{N}$ symmetric model
with periodic identifications $Z_{i+L}\equiv Z_{i}$ and $X_{i+L}\equiv X_{i}$. This model has a global symmetry generated by $\eta=\prod_{i}X_{i}$. Note that $\eta^{2}$,…, $\eta^{N-1}$ are also global symmetries. The model is known to have a critical point describing phase transition at $\lambda=1$ for $N=2,3,4$  [35, 36, 37]. At $\lambda=1$, there is a symmetry under the Kramers-Wannier transformation that interchanges the interaction term and the transverse-field term. This self-dual point is described by the $\frac{SU(2)_{N}}{U(1)}$ coset conformal field theory with central charge $c=\frac{2(N-1)}{N+2}$ [38, 39] for $N=2,3,4$. (We comment that for $N>4$, the central charge is $c=1$ [40, 41] and there is, in fact, an intermediate critical phase instead of a single transition point.) Let us denote the Kramers-Wannier duality symmetry for this model as $\mathbf{D}_{(N)}$. We define
$\begin{split}\mathbf{D}_{(N)}&\equiv\mathbf{P}_{(N)}\tilde{\mathbf{D}}_{(N)}\mathbf{P}_{(N)}\equiv\mathbf{P}_{(N)}\prod_{j=1}^{L-1}\left[\left(\sum_{n=0}^{N-1}c(n)X_{j}^{n}\right)\left(\sum_{n=0}^{N-1}\tilde{c}(n)(Z_{j}Z_{j+1}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}})^{n}\right)\right]\times\left(\sum_{n=0}^{N-1}c(n)X_{L}^{n}\right)\mathbf{P}_{(N)},\end{split}$ (11)
@@
$\tilde{c}(n)=\frac{\omega^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\frac{n(N+n)}{2}}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{N}}},\quad c(n)=\frac{\omega^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}\frac{n(N-n)}{2}}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{N}}}.$ (13)
@@
$\mathbf{D}_{(N)}X_{j}=Z_{j}Z_{j+1}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}\mathbf{D}_{(N)},$ (14a)
$\mathbf{D}_{(N)}X_{j}^{\dagger}=Z_{j}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}Z_{j+1}\mathbf{D}_{(N)},$ (14b)
$\mathbf{D}_{(N)}Z_{j}Z_{j+1}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}=X_{j+1}\mathbf{D}_{(N)},$ (14c)
$\mathbf{D}_{(N)}Z_{j}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}Z_{j+1}=X_{j+1}^{\dagger}\mathbf{D}_{(N)}.$ (14d)
Moreover, $\mathbf{D}_{(N)}$ commutes with the Hamiltonian at $\lambda=1$, $\mathbf{D}_{(N)}^{\dagger}\mathbf{D}_{(N)}=\mathbf{D}_{(N)}\mathbf{D}_{(N)}^{\dagger}=\mathbf{P}_{(N)}$ and
$\mathbf{D}_{(N)}^{2}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}e^{2i\phi(N,L)}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{T}_{clock}}\mathbf{P}_{(N)}\,.$ (15)
Here ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{T}_{clock}}$ is the translation on the lattice by one site and $e^{i\phi(N,L)}$ is a phase factor. For small values of $N$, we give the values of phase $e^{i\phi(N,L)}$ in Table. 1.
$N$ 2 3 4 5
$e^{i\phi(N,L)}$ $e^{\frac{2\pi iL}{8}}$ $e^{-\frac{2\pi iL}{4}}$ $e^{-\frac{2\pi iL}{8}}$ 1
Table 1: The values of phase $e^{i\phi(N,L)}$ for various choices of $N$.
@@
$\tilde{\mathbf{D}}_{(N)}Z_{i}=\omega^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{(N^{2}+1)}{2}}}Z_{1}\prod_{j=1}^{i}X_{j}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dagger}}\tilde{\mathbf{D}}_{(N)},$ (16)
which reduces to Eq. (6) for $N=2$.